OpenStudy (anonymous):

h(t)=t+2 g(t)=2t^2+3 Find h(-7) +g(-7)

OpenStudy (swag):

-7(h+g)

OpenStudy (anonymous):

nah its not like that....

OpenStudy (anonymous):

substitute t=-7 into both of h(t) and g(t) then do h(t=-7)+g(t=-7)

OpenStudy (anonymous):

h(-7)=-5, g(-7)=2(49)+3 h(-7)+g(-7)=-5+101=96

OpenStudy (anonymous):

yeah .. i know that.. but would that equation look like this if i were supossed to do that? h(t)=t+2 g(t)=2t^2+3 Find (h+g)(-7) <------

OpenStudy (anonymous):

yes cause all dependent variables are depend on only one independent variables, t so you can sum both of function together, what if h(-3) and g(-5) could you ??